Show that in Young's experiment the distance between successive maxima and minima are equal.


1
Expert's answer
2021-02-08T18:40:03-0500

The condition for constructive interference:


"d\text{ sin}\theta=m\lambda."


The distance between maxima for two successive m:


"\delta=(m+1)\lambda-m\lambda=\lambda."

The condition for destructive interference:


"d\text{ sin}\theta=\bigg(m+\frac12\bigg)\lambda."

The distance between two successive minima:


"\delta=\bigg(m+\frac12+1\bigg)\lambda-\bigg(m+\frac12\bigg)\lambda=\lambda."